How many water molecules are in a bottle of water that contains 500. mL? The density of water is 1.0 g/mL .
1
Expert's answer
2020-12-03T14:04:42-0500

"m= D*V"

1) 500*1.0=500 g

"n= m\/M"

2) n(H2O)= 500/(16+2)= 27,8 mole

"N=n*Na"

Na= 6,02*10^(23)

3) N= 27,8*6.02*10^(23) = 1.67356e+25